# Towards Flexible and Efficient Diffusion Low Light Enhancer

### v1

@@ Abstract
Diffusion-based Low-Light Image Enhancement (LLIE) has demonstrated significant success in improving the visibility of low-light images. However, the substantial computational burden introduced by the iterative sampling process remains a major concern. Current acceleration methods, whether training-based or training-free, often lead to significant performance degradation. As a result, to achieve an efficient student model with performance comparable to that of existing multi-step teacher model, it is usually necessary to retrain a more capable teacher model. This approach introduces inflexibility, as it requires additional training to enhance the teacher’s performance. To address these challenges, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), a step distillation framework specifically designed for LLIE. ReDDiT trains a student model to replicate the teacher’s trajectory in fewer steps while also possessing the ability to surpass the teacher’s performance. Specifically, we first introduce a trajectory decoder from the teacher model to provide guidance. Subsequently, a reflectance-aware trajectory refinement module is incorporated into the distillation process to enable more deterministic guidance from the teacher model. Our framework achieves comparable performance to previous diffusion-based methods with redundant steps in just 2 steps while establishing new state-of-the-art (SOTA) results with 8 or 4 steps. Comprehensive experimental evaluations on 10 benchmark datasets validate the effectiveness of our method, consistently outperforming existing SOTA methods.
Introduction
Low-light Image Enhancement (LLIE) is designed to improve the visibility and contrast of images captured in low-light conditions while maintaining natural-looking details, which is crucial for various downstream applications. Diffusion models have achieved remarkable success in this domain, demonstrating significant progress in generating photorealistic normal light images (Yi et al. 2023; Yin et al. 2023; Hou et al. 2024).
@@ Introduction
Previous diffusion-based LLIE methods primarily focus on how to condition the low-light image within the diffusion framework. Diff-Retinex applies the Retinex decomposition to the low-light image and condition these components into the diffusion model. (Yi et al. 2023). PyDiff utilizes pyramid low-light images as conditions and applies DDIM to expedite the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020). In WCDM, the diffusion model operates in the wavelet-transformed space, restoring the high-frequency parts of low-light images (Jiang et al. 2023). GSAD proposes a global structural regularization to enhance structural information learning, enabling smaller noise schedules for the inference stage (Hou et al. 2024). These methods predominantly focus on the conditioning of the LLIE and rely on universal acceleration techniques, i.e., the DDIM sampler, during the sampling,which results in poor trade-off between of the performance and efficiency.
We attempt to develop an efficient and flexible diffusion-based low-light enhancer with acceleration methods. However, a significant gap remains. Accelerating methods for diffusion models can generally be categorized into two types: training-based and training-free methods. Training-free methods, such as various ODE and SDE samplers, offer greater flexibility and convenience for compressing the sampling process. However, these methods typically maintain good performance only with 10 to 25 sampling steps. Training-based methods, including various distillation techniques (Salimans and Ho 2021; Song et al. 2023) and Rectified Flow methods (Liu 2022), can reduce the number of sampling steps to 10 through additional training. Unfortunately, these methods still suffer from significant performance degradation compared to the original model during compression, as highlighted in previous studies. This limitation implies that achieving an efficient diffusion model with comparable performance requires training a new teacher with enhanced capabilities, resulting in inflexibility.
Intuitively, if the student model can surpass the teacher during distillation, it could provide a flexible strategy for achieving efficient diffusion with strong performance. Recent distillation techniques face challenges as they operate on Gaussian noise, which introduces a larger search space. This approach leads to inefficiencies in fitting the teacher model’s trajectory, making it nearly impossible to achieve performance that exceeds the teacher model. While this issue persists in unconditional image generation, it can be addressed in LLIE since the degraded image is already available. This allows for the reduction of the search space by incorporating deterministic elements into the diffusion trajectory. The reflectance component of images, which captures the intrinsic characteristics of a scene and remains stable across varying illumination conditions, is a natural candidate for use as a deterministic prior to adjust and refine diffusion trajectories.
Motivated by this, we propose Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT). Firstly, We introduce a trajectory decoder for the teacher model to facilitate trajectory learning. Subsequently, recognizing the beneficial impact of the reflectance in low-light images, we propose a novel reflectance-aware trajectory refinement module. By incorporating illumination and noise level estimation, we predict the reflectance of low-light images based on the Retinex theory. Then the reflectance is remapped to the trajectory of the teacher diffusion model, refining its trajectory accordingly. Following this, trajectory matching distillation is conducted between the teacher and student models. To further improve the performance, auxiliary losses that directly obtain supervision signals from the data distribution are also integrated into the distillation process. Our method yields a 2-step diffusion model with comparable performance to previous 10-step diffusion-based methods, as well as 4 and 8-step diffusion models that achieve new SOTA results. In summary, our contributions are summarized as follows:
@@
(1) We introduce ReDDiT, a novel distillation scheme to enhance the efficiency of generative diffusion models for LLIE. Notably, ReDDiT achieves high-quality image restoration in just 2 steps, setting a new standard for efficient diffusion-based models in this domain.
(2) We design a Reflectance-Aware Trajectory Refinement (RATR) module, which utilizes the reflectance component of the low-light images to refine the overall trajectory of the teacher model. This refinement process incorporate deterministic guidance to the student model, enhancing its learning effectiveness.
@@
Related Work
Low-Light Image Enhancement
Enhancing low-light images is a classical task in low-level vision, with numerous solutions leveraging deep neural networks(Guo et al. 2020; Jiang et al. 2021; Wu et al. 2022). Recently, as the diffusion model has demonstrated promising results in image generation tasks, there has been growing interest in leveraging diffusion models for LLIE to achieve faithful restoration results. Diff-Retinex applies Retinex decomposition as the condition of diffusion model(Yi et al. 2023). PyDiff utilizes pyramid low-light images as the condition and applies DDIM to speed up the sampling stage (Zhou, Yang, and Yang 2023; Song, Meng, and Ermon 2020). In WCDM, the diffusion model is constructed in the wavelet-transformed space and restores the high-frequency parts of low-light images (Jiang et al. 2023). GSAD proposes a global structural regularization to enhance structural information learning, which also helps to train diffusion models with less curvature, enabling smaller noise schedules for the inference stage (Hou et al. 2024). Unfortunately, these works only consider how to condition the low-light images, neglecting the sampling strategy and raising efficiency concerns.
Diffusion Model Acceleration
Diffusion models have shown promising performance across various tasks such as style transfer (Chen et al. 2023), video generation (Deng et al. 2023), and text-to-image generation (Li et al. 2024; Ruiz et al. 2023). However, the issue of redundant sampling steps remains a significant efficiency concern. Recent research has explored various strategies to accelerate diffusion models. One series of methods involves the use of fast post-hoc samplers (Lu et al. 2022a, b; Song, Meng, and Ermon 2020), which reduces the number of inference steps for pre-trained diffusion models to 20-50 steps. However, most of them suffer from severe performance degradation when further accelerate sampling within 10 steps. To address this limitation, step distillation is proposed, aiming to distill diffusion models into fewer steps (Luhman and Luhman 2021; Salimans and Ho 2021). Progressive distillation (PD) (Salimans and Ho 2021) is the first successful practice and produces 2 step unconditional diffusion. Following PD, Meng et al. apply PD to the large-scale Stable Diffusion (Meng et al. 2023), achieving text-to-image generation with 2 steps. Consistency Distillation (CD) (Song et al. 2023) aims to learn consistency among diffusion timesteps of the teacher models (Song et al. 2023). Following this line of research, CTM extends such consistency from individual timesteps to an entire trajectory along the diffusion model, enabling faithful unconditional generation(Kim et al. 2023). Despite the booming development of diffusion distillation in other fields, techniques for LLIE are still left blank. In this paper, we present the first distilled diffusion tailored for LLIE, which will be introduced in detail in the following sections.
@@
Reflectance-Aware Diffusion Distillation
To address the challenges encountered in previous diffusion-based LLIE models, we propose ReDDiT, aiming to train a student diffusion model with parameters $\theta$ to learn the trajectory of the teacher model $\eta$ in fewer steps. In this section, we first introduce a trajectory decoder to represent the trajectory from the teacher model. Next, to incorporate the deterministic prior, we design a Reflectance-Aware Trajectory Refinement Module to refine the trajectory of the teacher model before performing distillation. In order to train a student model that surpasses the teacher model, we integrate auxiliary losses that receive direct supervision signals from the clean images during distillation. These aspects will be detailed below.
Trajectory Distillation
@@
The essence of diffusion distillation differs from traditional knowledge distillation is that distilled diffusion models learn the trajectory from the teacher, preserving the iterative sampling characteristics unique to diffusion models. We define $G(x_{t},y,t,s)$ as the trajectory decoder that transition from time step $t$ to $s$ $(s<t)$ to represent the diffusion model trajectory:
[…]
where $u\in[s,t)$ determines intermediate steps. Draw inspiration from the denoising diffusion implicit model, we represent the decoder with the teacher neural networks $\epsilon_{\eta}$ as:
$G_{\eta}(x_{t},y,t,s)$ $=\bar{\alpha}_{s}\epsilon_{\eta}(x_{t},y,t)+\sigma_{s}\frac{x_{t}-\bar{\alpha}_{t}\epsilon_{\eta}(x_{t},y,t)}{\sigma_{t}}$ (2)
$=\frac{\sigma_{s}}{\sigma_{t}}x_{t}+(\bar{\alpha}_{s}-\frac{\sigma_{s}}{\sigma_{t}}\bar{\alpha}_{t})\epsilon_{\eta}(x_{t},y,t),$
@@
with this, we can decode the trajectory of the teacher model from time step $s$ to $t$. For a more precise estimation of the entire trajectory, we also utilize the intermediate step $u\in[s,t)$ to estimate the trajectory, which formulates the trajectory decoder in the second order as:
$x^{\eta}_{s,u,t}$ $=G_{\eta}(G_{\eta}(x_{t},y,t,u),y,u,s),$ (3)
@@
with a perfectly pretrained teacher with parameters $\eta$, the trajectory decoder can accurately construct the trajectory. To facilitate trajectory learning during distillation, the student model should match the trajectory decoder from $t$ to $s$. The student model then learns to match this trajectory during distillation, represented as $x^{\theta}_{s,t}=G_{\eta}(x_{t},y,t,s)$, the distillation regularization is formulated as:
$\mathcal{L}_{distill}=\|x^{\theta}_{s,t}-x^{\eta}_{s,u,t}\|_{2}^{2},$ (4)
@@
with such regularization, the information intermediate step $u\in[s,t)$ will be distilled in student model. In practice, we utilize the clean image predicted by the trajectory both on the student and teacher to perform distillation for stable training. The predicted clean images of the teacher and student denoted as $x_{target}$ and $x_{est}$, are formulated as follows:
@@
where $\lambda(t)$ is an adaptive weight that controls the distillation according to the importance schedule. Drawing inspiration from the vanilla DDPM (Ho, Jain, and Abbeel 2020), the $\lambda(t)$ is set as $\omega(t)=\max(1,\frac{\bar{\alpha}_{t}^{2}}{\sigma_{t}^{2}})$.

### v2

# Efficient Diffusion as Low Light Enhancer

@@ Abstract
The computational burden of the iterative sampling process remains a major challenge in diffusion-based Low-Light Image Enhancement (LLIE). Current acceleration methods, whether training-based or training-free, often lead to significant performance degradation, highlighting the trade-off between performance and efficiency. In this paper, we identify two primary factors contributing to performance degradation: fitting errors and the inference gap. Our key insight is that fitting errors can be mitigated by linearly extrapolating the incorrect score functions, while the inference gap can be reduced by shifting the Gaussian flow to a reflectance-aware residual space. Based on the above insights, we design Reflectance-Aware Trajectory Refinement (RATR) module, a simple yet effective module to refine the teacher trajectory using the reflectance component of images. Following this, we introduce Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), an efficient and flexible distillation framework tailored for LLIE. Our framework achieves comparable performance to previous diffusion-based methods with redundant steps in just 2 steps while establishing new state-of-the-art (SOTA) results with 8 or 4 steps. Comprehensive experimental evaluations on 10 benchmark datasets validate the effectiveness of our method, consistently outperforming existing SOTA methods.
[…]
1 Introduction
Figure 1: ReDDiT shifts the teacher trajectory from the original Gaussian distribution to a residual space, effectively reducing the sampling gap. Subsequently, it refines the teacher trajectory towards the ground truth trajectory to mitigate fitting errors.
@@ 1 Introduction
Low-light Image Enhancement (LLIE) is designed to improve the visibility and contrast of images captured in low-light conditions while maintaining natural-looking details, which is crucial for various downstream applications. Diffusion models [9, 22] have achieved remarkable success in this domain, demonstrating significant progress in generating photorealistic normal light images [40, 41, 10]. Previous diffusion-based LLIE methods [40, 41, 10] primarily focus on how to condition the low-light image within the diffusion framework. Diff-Retinex [40] applies the Retinex decomposition to the low-light image and condition these components into the diffusion model. PyDiff [48] utilizes pyramid low-light images as conditions and applies DDIM [26] to expedite the sampling stage. In WCDM [11], the diffusion model operates in the wavelet-transformed space, restoring the high-frequency parts of low-light images. GSAD [10] proposes a global structural regularization to enhance structural information learning, enabling smaller noise schedules for the inference stage.
However, the primary challenge to the broader application of diffusion models in LLIE stems from their iterative denoising mechanism. For example, DDPM [9] requires multiple denoising steps, e.g., 1000 steps, to reverse a Gaussian noise into a clean image. The computation overhead conflicts with the demands of LLIE, particularly in computational photography applications for edge devices like mobile phones and surveillance cameras.
Inspired by training-aware acceleration methods for diffusion models [25, 27], we investigate distilling a multi-step diffusion-based LLIE model to improve efficiency. A key observation is that performance degradation is inevitable as the number of sampling steps is reduced, even when using advanced acceleration methods (e.g., consistency distillation [27]). This leads us to pose the question: Is it possible to distill a student diffusion model that surpasses the original diffusion model? If so, even as performance degrades with fewer sampling steps, we could still achieve results comparable to those of the multi-step teacher model.
To address this, we conduct a comprehensive analysis of diffusion model acceleration techniques. Through theoretical analysis, we identify two primary factors contributing to performance degradation: (1) The fitting error. It is the unavoidable errors between deep learning models and target fitting data. It will result the additional undesired terms and mismatch during distillation. (2) The inference gap. It is the gap between the training target and sampling strategies specialized for the Diffusion models. It is caused by the universal diffusion model often trained and operated on the Gaussian flow for the diversity of generation, while the LLIE requires for more deterministic.
Our key insight is to refine the teacher trajectory by applying linear extrapolation to the score functions of the teacher model, thereby mitigating the adverse effects of fitting errors. Meanwhile, shifting the sampling trajectory to a deterministic space addresses the sampling gap, as shown in Figure 1. A detailed theoretical analysis of this approach is presented in  Sec. 3.
Based on the above principles, we design Reflectance-Aware Trajectory Refinement (RATR) module, a simple yet effective module to refine the teacher trajectory for LLIE task. It incorporates the reflectance component of images as a deterministic prior to adjust and refine diffusion trajectories. Following this, we introduce Reflectance-aware Diffusion with Distilled Trajectory (ReDDiT), an efficient and flexible distillation framework tailored for LLIE. This framework perform trajectory matching distillation is conducted between the refined teacher and student models, yields a 2-step diffusion model with comparable performance to previous 10-step diffusion-based methods, as well as 4 and 8-step diffusion models that achieve new SOTA results. our contributions are summarized as follows:
(1) We theoretically analyze the factors contributing to performance degradation and propose targeted improvements: linear extrapolation of the score function to mitigate performance loss due to fitting errors, and residual shifting as a solution to the sampling gap.
@@
(2) Based on these two design principles, we present ReDDiT, a novel distillation scheme that enhances the efficiency of generative diffusion models for LLIE. Notably, ReDDiT achieves high-quality image restoration in just 2 steps, establishing a new benchmark for efficient diffusion-based models in this domain.
@@
2 Related Work
Low-Light Image Enhancement. Enhancing low-light images is a classical task in low-level vision, with numerous solutions leveraging deep neural networks [7, 12, 36]. Recently, as the diffusion model has demonstrated promising results in image generation tasks, there has been growing interest in leveraging diffusion models for LLIE to achieve faithful restoration results. Diff-Retinex applies Retinex decomposition as the condition of diffusion model [40]. PyDiff utilizes pyramid low-light images as the condition and applies DDIM to speed up the sampling stage [48, 26]. In WCDM, the diffusion model is constructed in the wavelet-transformed space and restores the high-frequency parts of low-light images [11]. GSAD proposes a global structural regularization to enhance structural information learning, which also helps to train diffusion models with less curvature, enabling smaller noise schedules for the inference stage [10]. Unfortunately, these works only consider how to condition the low-light images, neglecting the efficiency concerns.
Diffusion Model Acceleration. Diffusion models have shown promising performance across various tasks such as style transfer [4], video generation [5], and text-to-image generation [16, 23, 22]. However, the issue of redundant sampling steps remains a significant efficiency concern. Recent research has explored various strategies to accelerate diffusion models. One series of methods involves the use of fast post-hoc samplers [17, 18, 26], which reduces the number of inference steps for pre-trained diffusion models to 20-50 steps. However, most suffer from severe performance degradation when further accelerating sampling within 10 steps. To address this limitation, step distillation [19, 25] is proposed, aiming to distill diffusion models into fewer steps. Progressive distillation (PD) [25] is the first successful practice and produces 2-step unconditional diffusion. Following PD [25], Meng et al. apply PD to the large-scale Stable Diffusion, achieving text-to-image generation with 2 steps. Consistency Distillation (CD) [27] aims to learn consistency among diffusion timesteps of the teacher models [27]. Following this line of research, CTM [13] extends such consistency from individual timesteps to an entire trajectory along the diffusion model, enabling faithful unconditional generation. Despite the booming development of diffusion distillation in other fields, techniques for LLIE are still left blank. In this paper, we present the first distilled diffusion tailored for LLIE, which will be introduced in detail in the following sections.
@@ 2 Related Work
3 Methods
Let $x_{0}$ denote a high-quality image, the forward diffusion process aims to generate a sequence of noisy latent variables $x_{1},x_{2},...,x_{T}$ using a Markovian process, which is defined by the equation:
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon,$ (1)
where $\alpha_{t}\in(0,1)$ represents the noise schedule, $\sigma_{t}$ denotes the covariance at $t$, and $\epsilon$ represents the Gaussian noise. In the reverse process, we incorporate the low-light image $y$ as the conditioning of the score functions $\epsilon_{\eta}(x_{t},y,\alpha_{t})$ and predict the clean images iteratively:
$x^{\eta}_{t-1}=\alpha_{t-1}\left(\frac{x_{t}-\sigma_{t}\epsilon_{\eta}}{\alpha_{t}}\right)+\sigma_{t-1}\epsilon_{\eta}.$ (2)
In this section, we propose ReDDiT, aiming to train a student diffusion model with parameters $\theta$ to learn the trajectory of the teacher model $\eta$ in fewer steps. We begin by introducing trajectory distillation. Next, we theoretically analyze the factors contributing to performance degradation and present a concise formulation for refining the trajectory. Then, we determine that the reflectance component is most effective for refining the teacher trajectory and design a Reflectance-Aware Trajectory Refinement (RATR) module. Finally, all components are integrated into the ReDDiT framework, where distillation is performed.
3.1 Trajectory Distillation
@@
The essence of diffusion distillation differs from traditional knowledge distillation is that distilled diffusion models learn the trajectory from the teacher, preserving the iterative sampling characteristics unique to diffusion models. We define $G(x_{t},y,t,s)$ as the trajectory decoder that transition from time step $t$ to $s$ $(s<t)$ to represent the diffusion model trajectory. Draw inspiration from the denoising diffusion implicit model, we represent the decoder with the student score function $\epsilon_{\theta}$ as:
$G_{\theta}(x_{t},y,t,s)=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t).$ (3)
@@
With this trajectory decoder, the teacher model’s trajectory can be decoded from time step $s$ to $t$. For a more precise estimation of the entire teacher trajectory, we also utilize the intermediate step $u\in[s,t)$ to estimate the trajectory, which formulates the trajectory decoder in the second order as:
$x^{\eta}_{s,u,t}$ $=G_{\eta}(G_{\eta}(x_{t},y,t,u),y,u,s)$ (4)
$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$
$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}(\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u))$
$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t),$
@@
To facilitate trajectory learning during distillation, the student model should match the teacher trajectory from $t$ to $s$. Denoting the student trajectory as $x^{\theta}_{s,t}=G_{\eta}(x_{t},y,t,s)$, the distillation regularization is formulated as:
$\mathcal{L}_{distill}=\|x^{\theta}_{s,t}-x^{\eta}_{s,u,t}\|_{2}^{2}.$ (5)
@@
With such regularization, the information intermediate step $u\in[s,t)$ will be distilled in student model. In practice, we utilize the clean image predicted by the trajectory both on the student and teacher to perform distillation for stable training. The predicted clean images of the teacher and student, denoted as $x_{target}$ and $x_{est}$, are formulated as follows:
@@
where $\lambda(t)$ is an adaptive weight and is set as $\max(1,\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}})$.
3.2 On the Refinement of the Teacher Trajectory
Direct application of trajectory distillation often leads to significant performance degradation. In this section, we theoretically analyze the core reasons behind this degradation and propose a strategy to mitigate its effects.
On the fitting error of the teacher trajectory. To understand this, let us revisit Equation 3 and Equation 4, and consider the ideal condition where $\mathcal{L}_{distill}=0$, meaning that the student’s trajectory perfectly matches the teacher’s trajectory. Under this condition, we have:
$(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t)$ $=\sigma_{s}\epsilon_{\eta}(x_{u},y,u)-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t)$ (8)
$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}.$
Since the teacher model is trained using the vanilla diffusion loss function, the ideal teacher trajectory satisfies the condition $\epsilon_{\eta}(x_{t},y,t)=\epsilon_{\eta}(x_{u},y,u)=\tilde{\epsilon}$. Under this condition, the Equation 8 can be simplified as:
$(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\epsilon_{\theta}(x_{t},y,t)$ $=(\sigma_{s}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t})\tilde{\epsilon}.$ (9)
However, the existence of fitting errors makes it impossible to achieve this goal. The presence of guidance with undesired components inevitably leads to suboptimal results.
On the mitigation of the fitting error. Fortunately, we find that these undesired terms can be mitigated by a scaling parameter $\omega\in(0,1]$. For the term $\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$, we can further reduce its impact by multiplying it by the scaling parameter $\omega$. The mismatch between $\epsilon_{\eta}(x_{u})$ and $\epsilon_{\eta}(x_{t})$ introduces a more significant optimization conflict. Due to the presence of fitting error, this term is not always consistent with $\epsilon_{\eta}(x_{t},y,t)$. Our key insight is to refine this term through linear interpolation toward the ideal value, expressed as $\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}$. This approach not only maintains consistency with previous terms but also provides linear trajectory guidance. After applying these operations, the refined teacher trajectory can be expressed as follows:
$\tilde{x}^{\eta}_{s,u,t}$ $=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$ (10)
$+\omega\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$
$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}.$
To this end, the distillation loss, $\mathcal{L}_{distill}$, is equivalent to the following, disregarding the coefficients of each term:
$\mathcal{L}_{distill}$ $=||\epsilon_{\theta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)||$ (11)
$+||\epsilon_{\theta}(x_{t},y,t)-(\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon})||.$
On the mitigation of the inference gap. Considering the vanilla diffusion are trained on the Gaussian flow, we determine the $\tilde{\epsilon}$ not the pure Gaussian noise but shift it into a residual space as:
$\tilde{\epsilon}=\frac{x_{t}-\alpha_{t}\tilde{x}_{0}}{\sigma_{t}}=\frac{\alpha_{t}(x_{0}-\tilde{x}_{0})}{\sigma_{t}}+\epsilon,$ (12)
where $\tilde{{x}}_{0}$ should lie between the low-light and clean image spaces, serving as an intermediate initial space for the student model to learn. This positioning ensures a closer initial distribution for the student model compared to the Gaussian distribution
Additionally, our investigation shows that the refinement of the teacher trajectory in Equation 10 can be implemented through a straightforward approach, as outlined in the Corollary 1.
Corollary 1 (Proof in the supplementary material).
Given the refinement component $\tilde{x}_{s}=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$, the Equation 10 is equivalent to : $\tilde{x}^{\eta}_{s,u,t}=\omega x^{\eta}_{s,u,t}+(1-\omega)\tilde{x}_{s}$.
3.3 Reflectance-Aware Trajectory Refinement
On the determination of the component $\tilde{x}_{0}$, using the ground truth clean images $x_{0}$ will cause $\tilde{\epsilon}$ degrade to $\epsilon$. A natural approach is to use the low-light images $y$ as $\tilde{x}_{0}$. In practice, the reflectance, which shares characteristics with both the clean images and the low-light images $y$, serves as a better component for refining the trajectory. Building on this observation, we propose the RATR module to refine the teacher’s trajectory, thereby reducing inference gap.
Given an illumination map $h$ and the noise $z$, the reflectance can be obtained through $x=\frac{y-z}{h}$ based on the Retinex theory. For illumination estimation, we employ the maximum channel of the low-light image $y$ to represent the estimated illumination map $h^{\prime}$, as the common practice. Regarding ISO noise estimation, similar to the previous non-learning-based denoising method [1], the noise can be modeled as the distance between the noisy image and the clean image. We use this distance to estimate the noise map of the input low-light images:
$z^{\prime}=|y-\psi(y)|,$ (13)
where $\psi$ represents a non-learning-based denoising operation, allowing for flexible distillation. With these estimations, we can obtain a latent clean image $\tilde{x}_{0}=\frac{y-z^{\prime}}{h^{\prime}}$. The $\tilde{x}_{0}$ is then employed to the trajectory refinement to perform distillation. Extensive experiments in ablation studies show that such refinement reduce the inference gap and greatly improves performance.